An aircraft at a distance of 500m from the control tower of an air strip begins to taxi his way on a runway ,from the tower, with a speed of 100km/hr. If one of the pilots in the cockpit begins to run towards the tail section of an aircraft with a speed of 1m/s, how fast will the observer in the control tower,see the pilot approaching the tail section?
1
Expert's answer
2018-09-06T16:34:08-0400
The speed of the pilot approaching the tail section:
V_p=V-v
V_p=100/3.6-1=26.8 m/s.